joey is thrice as old as his son. if the sum of their ages is 60. how old is each of them?
Answer by mananth(16946)

let son's age =x
Joey's age = 3x
their sum = 60
x+3x=60
4x=60
/4
x= 15
son's age = 15
Joey's age = 15*3 = 45